What is the concentration of ClO– in a 0.010 M solution of HClO at 25°C? The...

  1. What is the concentration of ClO– in a 0.010 M solution of HClO at 25°C? The value of Ka at this temperature is 3.0 × 10−8.
Answer #1

                                                                  HClO + H2O H3O+ + ClO-

Initial concentration (M):                            0.010                  0            0

Change in concentration (M):                     -X                      X            X

Equilibrium concentration (M):                (0.010 - X)            X             X

Ka of HClO = 3.0 × 10−8

Now,

Ka = [H3O+][ClO-]/[HClO]

or, 3.0 × 10−8 = (X)(X)/(0.010 - X)

or, 3.0 × 10−8 = X2/0.010

or, X2 = 0.0000000003

or, X = 1.73 x 10-5

or, [ClO-] = 1.73 x 10-5 M

Therefore, the concentration of ClO- in a 0.010 M solution of HClO = 1.73 x 10-5 M